OpenStudy (anonymous):

Find the slope and y intercept for the given line. 3 x + 2 y = 7 m = y -intercept Find the slope and y intercept for the given line. 3 x + 2 y = 7 m = y -intercept @Mathematics

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

to solve this, you have to get the line (which is in standard form right now) into slope-intercept form (y=mx+b; m is the slope, b is the y-intercept.) get y on a side by itself. you will have to subtract 3x from the left side and the right side. 3x+2y=7 -3x -3x ------------- 2y=7-3x now divide the 2 from the left side and the right side to get y by itself. 2y/2y=7-3x/2y y=-3x+7 thus, your slope (m) is -3 and your y intercept (b) is 7.

OpenStudy (anonymous):

sorry, i did my division incorrectly. when you divide 2 from both sides you get y=-3/2x-7/2. therefore your slope is -3/2 and your y intercept is 7/2

OpenStudy (anonymous):

ChrisS Can you help me?

OpenStudy (anonymous):

Solve by using a two-order system. Ben has $3.40 consisting of quarters and dimes. How many coins of each kind does he have if he has 22 coins? _____quarters and_____ dimes